# Policy-Driven World Model Adaptation for Robust Offline Model-based Reinforcement Learning

### v2

@@ 4 Policy-guided World Model Adaptation for Enhanced Robustness
To obtain the theoretical guarantee shown in Section 3, the maximin problem in Eq. (5) needs to be well solved. A straightforward approach, as employed in [DBLP:conf/nips/RigterLH22], is to update the policy $\pi_{\theta}$ and world model $P_{\phi}$ alternatively, treating the other as part of the environment during each update step. Specifically, the training process at iteration $k$ is as follows:
$\theta^{k+1}=\theta^{k}+\eta_{\theta}^{k}\nabla_{\theta}J(\theta^{k},\phi^{k});\ \phi^{k+1}=\phi^{k}-\eta_{\phi}^{k}\nabla_{\phi}\mathcal{L}(\theta^{k},\phi^{k}).$ (7)
Here, $\mathcal{L}(\theta^{k},\phi^{k})=J(\theta^{k},\phi^{k})+\lambda\mathbb{E}_{(s,a)\sim\mathcal{D}_{\mu}}\left[KL(P_{\bar{\phi}}(\cdot|s,a)||P_{\phi^{k}}(\cdot|s,a))-\epsilon\right]$, $\eta_{\theta}^{k},\eta_{\phi}^{k}$ are learning rates at iteration $k$. In this case, the constrained optimization problem $\min_{\phi\in\Phi}J(\theta,\phi)$ is relaxed into an unconstrained optimization problem $\min_{\phi\in\mathcal{M}}\mathcal{L}(\theta,\phi)$, by employing the Lagrangian formulation and treating the Lagrange multiplier $\lambda>0$ as a hyperparameter. This method is simple but lacks formal convergence guarantees. The alternating updates may lead to instability due to the non-stationarity introduced by treating the other model as part of the environment in separate updates.
By viewing the policy and world model as the leader and follower in a general-sum Stackelberg game, we can apply the Stackelberg learning dynamics (i.e., Eq. (4)) to iteratively update the policy and world model. Such learning dynamics have proven effective in robust online RL [DBLP:conf/ijcai/HuangXFZ22] and offline model-free RL [DBLP:journals/corr/abs-2309-16188]. In particular, the update at iteration $k$ is as follows:
$\theta^{k+1}=\theta^{k}+\eta_{\theta}^{k}\left[\nabla_{\theta}J(\theta^{k},\phi^{k})-\nabla^{2}_{\phi\theta}\mathcal{L}(\theta^{k},\phi^{k})^{T}(\nabla^{2}_{\phi}\mathcal{L}(\theta^{k},\phi^{k}))^{-1}\nabla_{\phi}J(\theta^{k},\phi^{k})\right];$ (8)
@@
The formula above can be derived by substituting $(x_{1},x_{2})$ and $(f_{1},f_{2})$ in Eq. (4) with $(\theta,\phi)$ and $(-J,\mathcal{L})$, respectively. Compared to Eq. (7), the policy updates incorporate gradient information from the world model, as the best-response world model is inherently a function of the policy. This dependency allows the policy optimization process to account for the influence of the evolving world model, leading to potentially more stable learning dynamics. As noted in Section 2, models updated using such learning dynamics are guaranteed to converge to a local Stackelberg equilibrium under mild conditions.
@@
The second approach still solves an unconstrained problem $\min_{\phi\in\mathcal{M}}\mathcal{L}(\theta,\phi)$. As an improvement, we propose a novel learning dynamics, where the follower $\theta$ solves the constrained problem $\min_{\phi\in\Phi}J(\theta,\phi)$ as a response to the leader:
@@
$\theta^{k+1}=\theta^{k}+\eta_{\theta}^{k}\left[\nabla_{\theta}J(\theta^{k},\phi^{k})-\nabla^{2}_{\phi\theta}\mathcal{L}(\theta^{k},\phi^{k},\lambda^{k})^{T}H(\theta^{k},\phi^{k},\lambda^{k})\nabla_{\phi}J(\theta^{k},\phi^{k})\right];$ (9)
$\phi^{k+1}=\phi^{k}-\eta_{\phi}^{k}\nabla_{\phi}\mathcal{L}(\theta^{k},\phi^{k},\lambda^{k}),\ \lambda^{k+1}=[\lambda^{k}+\eta_{\lambda}^{k}\nabla_{\lambda}\mathcal{L}(\theta^{k},\phi^{k},\lambda^{k})]^{+}.$
@@
In the equation above, $H(\theta^{k},\phi^{k},\lambda^{k})=A^{-1}+\lambda^{k}A^{-1}BS^{-1}B^{T}A^{-1}$, $S=C-\lambda^{k}B^{T}A^{-1}B$, $A=\nabla^{2}_{\phi}\mathcal{L}(\theta^{k},\phi^{k},\lambda^{k})$, $B=\nabla^{2}_{\phi\lambda}\mathcal{L}(\theta^{k},\phi^{k},\lambda^{k})$, $C=\nabla_{\lambda}\mathcal{L}(\theta^{k},\phi^{k},\lambda^{k})$, and $\mathcal{L}(\theta^{k},\phi^{k},\lambda^{k})=J(\theta^{k},\phi^{k})+\lambda^{k}\mathbb{E}_{(s,a)\sim\mathcal{D}_{\mu}}\left[KL(P_{\bar{\phi}}(\cdot|s,a)||P_{\phi^{k}}(\cdot|s,a))-\epsilon\right]$ is the Lagrangian. Eq. (9) presents how to update the dual variable in the constrained optimization problem $\min_{\phi\in\Phi}J(\theta,\phi)$, i.e., $\lambda$, alongside $\theta$ and $\phi$. Compared to Eq. (8), the policy updates additionally incorporate gradient information from the best-response dual variable, which is a function of $\pi_{\theta}$. This allows the policy optimization to account for the implicit influence of constraint satisfaction, leading to a more informed learning process. The second line in Eq. (9) corresponds to a primal-dual method for solving $\min_{\phi\in\Phi}J(\theta,\phi)$, where $\cdot$$^{+}$ is a projection operation onto the non-negative real space. Widely used in constrained RL [^4] Viewing $P_{\phi}$ and $\pi_{\theta}$ as the ”policy” and ”world model” respectively, $\min_{\phi\in\Phi}J(\theta,\phi)$ is converted into a typical constrained RL problem., this method can achieve low suboptimality if the world model’s parameterization has sufficient representational capacity, as detailed in [DBLP:journals/tac/PaternainCCR23]. Additionally, the first line in Eq. (9) represents a gradient ascent step corresponding to the objective function $\max_{\theta}\{J(\theta,\phi^{*}(\theta))\mid\phi^{*}(\theta)\in\arg\min_{\phi\in\Phi}J(\theta,\phi)\}$, of which the derivation is provided in Appendix F. According to the learning dynamics in Stackelberg games and constrained optimization [DBLP:conf/icml/FiezCR20, DBLP:journals/tac/PaternainCCR23], the learning rates in Eq. (9) should satisfy $\eta_{\phi}^{k}\gg\eta_{\lambda}^{k}\gg\eta_{\theta}^{k}$ to ensure convergence.
@@ 5 Practical Algorithm: ROMBRL
For a discounted finite-horizon MDP with horizon $h$, $J(\theta,\phi)=\mathbb{E}_{\tau\sim P(\cdot;\theta,\phi)}\left[\sum_{j=0}^{h-1}\gamma^{j}r_{j}\right]$, where $\tau=(s_{0},a_{0},r_{0},\cdots,s_{h})$ and $P(\tau;\theta,\phi)=d_{0}(s_{0})\prod_{j=0}^{h-1}\pi_{\theta}(a_{j}|s_{j})P_{\phi}(r_{j},s_{j+1}|s_{j},a_{j})$. To compute the first-order and second-order derivatives in Eqs. (7) - (9), we have the following theorem:
Theorem 4.
@@
$\nabla_{\theta}J(\theta,\phi)=\mathbb{E}_{\tau\sim P(\cdot;\theta,\phi)}\left[\nabla_{\theta}\Psi(\tau,\theta)\right],\ \nabla_{\phi}J(\theta,\phi)=\mathbb{E}_{\tau\sim P(\cdot;\theta,\phi)}\left[\nabla_{\phi}\Psi(\tau,\phi)\right],$ (10)
@@
$\nabla_{\phi}^{2}J(\theta,\phi)=\mathbb{E}_{\tau\sim P(\cdot;\theta,\phi)}\left[\nabla_{\phi}\Psi(\tau,\phi)\nabla_{\phi}\log P(\tau;\theta,\phi)^{T}+\nabla_{\phi}^{2}\Psi(\tau,\phi)\right].$
@@
In a practical algorithm, the expectation terms in Eqs. (10) and (54) are replaced with corresponding unbiased estimators, i.e., the sample means. Estimators of the first-order derivatives can be efficiently computed using automatic differentiation, whose space and time complexity scale linearly with the number of parameters in the models, namely $N_{\theta}$ and $N_{\phi}$. However, computing the second-order derivatives, i.e., $\nabla^{2}_{\phi}\Psi(\pi,\phi)$ in $\nabla^{2}_{\phi}J(\theta,\phi)$ and $\nabla^{2}_{\phi}\log P_{\phi}$ in $\nabla^{2}_{\phi}\mathcal{L}(\theta,\phi,\lambda)$, can be costly. Instead of substituting the second-order terms with $cI$ as in [wang2021learning], where $I$ is the identity matrix, we propose the following approximations:
$\mathbb{E}_{\tau\sim P(\cdot;\theta,\phi)}\left[\nabla_{\phi}^{2}\Psi(\tau,\phi)\right]\approx-\mathbb{E}_{\tau\sim P(\cdot;\theta,\phi)}\left[\sum_{i=0}^{h-1}\left(\sum_{j=i}^{h-1}\gamma^{j}r_{j}\right)F(s_{i},a_{i},r_{i},s_{i+1};\phi)\right],$ (11)
$\mathbb{E}_{(s,a,r,s^{\prime})\sim P_{\bar{\phi}}\circ\mathcal{D}_{\mu}(\cdot)}\left[\nabla^{2}_{\phi}\log P_{\phi}(r,s^{\prime}|s,a)\right]\approx-\mathbb{E}_{(s,a,r,s^{\prime})\sim P_{\bar{\phi}}\circ\mathcal{D}_{\mu}(\cdot)}\left[F(s,a,r,s^{\prime};\phi)\right],$
where $F(s,a,r,s^{\prime};\phi)=\nabla_{\phi}\log P_{\phi}(r,s^{\prime}|s,a)\nabla_{\phi}\log P_{\phi}(r,s^{\prime}|s,a)^{T}$ is the Fisher Information Matrix. Thus, we substitute $\nabla^{2}_{\phi}\log P_{\phi}$ with F-F, inspired by the fact [amari2016information] that $\mathbb{E}_{x\sim P_{\phi}(\cdot)}\left[\nabla^{2}_{\phi}\log P_{\phi}(x)+\nabla_{\phi}\log P_{\phi}(x)\nabla_{\phi}\log P_{\phi}(x)^{T}\right]=0$. In Appendix H, we discuss the approximation errors that arise when using Eq. (11).
The other computational bottleneck is computing the inverse matrix $A^{-1}$ in Eq. (9) [^5] Since $S$ in Eq. (9) is a scalar, its inverse $S^{-1}$ can be easily computed.. We choose not to use an identity matrix in place of $A^{-1}$ as done in [DBLP:conf/aaai/NikishinAAB22]. As aforementioned, $A=\nabla^{2}_{\phi}\mathcal{L}(\theta,\phi,\lambda)$ is substituted with its sample-based estimator $\hat{A}$, which is defined as follows:
@@
where $U,V\in\mathbb{R}^{N_{\phi}\times m}$; $X,Y,Z\in\mathbb{R}^{N_{\phi}\times M}$; $m$ and $M$ represent the number of sampled trajectories and sampled transitions, respectively. Specifically, the $i$-th columns of $U$ and $V$ are given by $\nabla_{\phi}\Psi(\tau(i),\phi)/\sqrt{m}$ and $\nabla_{\phi}\log P(\tau(i);\theta,\phi)/\sqrt{m}$, respectively, where $\tau(i)$ denotes the $i$-th trajectory sampled from $P(\cdot;\theta,\phi)$. The $i$-th column of $Z$ is given by $\nabla_{\phi}\log P_{\phi}(r_{i},s^{\prime}_{i}|s_{i},a_{i})\cdot\sqrt{\lambda/M}$, based on a transition sampled from $P_{\bar{\phi}}\circ\mathcal{D}_{\mu}(\cdot)$. Additionally, by randomly sampling a trajectory $\tau$ and a time step $t\sim\text{Uniform}(0,h-1)$, we obtain the columns of $X$ and $Y$ as $\left(\sum_{j=t}^{h-1}\gamma^{j}r_{j}\right)\nabla_{\phi}\log P_{\phi}(s_{t+1},r_{t}|s_{t},a_{t})\cdot\sqrt{h/M}$ and $\nabla_{\phi}\log P_{\phi}(s_{t+1},r_{t}|s_{t},a_{t})\cdot\sqrt{h/M},$ respectively. Given that $m,M\ll N_{\theta},N_{\phi}$, each term in $\hat{A}$ is a low rank matrix and so we can apply Woodbury matrix identity [DBLP:conf/icml/FiezCR20] to efficiently compute $\hat{A}^{-1}$.

### v3

@@ 4 Policy-guided World Model Adaptation for Enhanced Robustness
To obtain the theoretical guarantee shown in Section 3, the maximin problem in Eq. (5) needs to be well solved. A straightforward approach, as employed in Rigter et al. (2022), is to update the policy $\pi_{\theta}$ and world model $P_{\phi}$ alternatively, treating the other as part of the environment during each update step. Specifically, the training process at iteration $k$ is as follows:
$\theta^{k+1}=\theta^{k}+\eta_{\theta}^{k}\nabla_{\theta}J(\theta^{k},\phi^{k});\$(7)
@@
Here, the objective function is defined as: $\mathcal{L}(\theta^{k},\phi^{k})=J(\theta^{k},\phi^{k})+\lambda\mathbb{E}_{\mathcal{D}_{\mu}}\left[KL(P_{\bar{\phi}}(\cdot|s,a)||P_{\phi^{k}}(\cdot|s,a))-\epsilon\right],$ where $\eta_{\theta}^{k},\eta_{\phi}^{k}$ are learning rates at iteration $k$. In this case, the constrained optimization problem $\min_{\phi\in\Phi}J(\theta,\phi)$ is relaxed into an unconstrained optimization problem $\min_{\phi\in\mathcal{M}}\mathcal{L}(\theta,\phi)$, by employing the Lagrangian formulation and treating the Lagrange multiplier $\lambda>0$ as a hyperparameter. This method is simple but lacks formal convergence guarantees. Alternating updates risk instability due to the non-stationarity introduced by treating the other model as part of the environment in separate updates.
By viewing the policy and world model as the leader and follower in a general-sum Stackelberg game, we can apply the Stackelberg learning dynamics (i.e., Eq. (4)) to iteratively update the policy and world model. Such learning dynamics have proven effective in robust online RL (Huang et al., 2022) and offline model-free RL (Zhou and Qu, 2023). In particular, the update at iteration $k$ is as follows:
$\theta^{k+1}$ $=\theta^{k}+\eta_{\theta}^{k}\bigg[\nabla_{\theta}J(\theta^{k},\phi^{k})$ (8)
$\quad-\nabla^{2}_{\phi\theta}\mathcal{L}(\theta^{k},\phi^{k})^{T}(\nabla^{2}_{\phi}\mathcal{L}(\theta^{k},\phi^{k}))^{-1}\nabla_{\phi}J(\theta^{k},\phi^{k})\bigg];$
@@
The formula above can be derived by substituting $(x_{1},x_{2})$ and $(f_{1},f_{2})$ in Eq. (4) with $(\theta,\phi)$ and $(-J,\mathcal{L})$, respectively. Compared to Eq. (7), policy updates integrate model gradients, as the best-response world model is inherently a function of the policy. This dependency allows the policy optimization process to account for the influence of the evolving world model, leading to potentially more stable learning dynamics. As noted in Section 2, models updated using such learning dynamics are guaranteed to converge to a local Stackelberg equilibrium under mild conditions.
@@
The second approach still solves an unconstrained problem $\min_{\phi\in\mathcal{M}}\mathcal{L}(\theta,\phi)$. As an improvement, we propose a novel learning dynamics, where the follower $\theta$ solves the constrained problem $\min_{\phi\in\Phi}J(\theta,\phi)$ responding to the leader:
$\theta^{k+1}$ $=\theta^{k}+\eta_{\theta}^{k}\bigg[\nabla_{\theta}J(\theta^{k},\phi^{k})$ (9)
$\quad-\nabla^{2}_{\phi\theta}\mathcal{L}(\theta^{k},\phi^{k},\lambda^{k})^{T}H(\theta^{k},\phi^{k},\lambda^{k})\nabla_{\phi}J(\theta^{k},\phi^{k})\bigg];$
$\phi^{k+1}$ $=\phi^{k}-\eta_{\phi}^{k}\nabla_{\phi}\mathcal{L}(\theta^{k},\phi^{k},\lambda^{k});$
$\lambda^{k+1}$ $=\big[\lambda^{k}+\eta_{\lambda}^{k}\nabla_{\lambda}\mathcal{L}(\theta^{k},\phi^{k},\lambda^{k})\big]^{+}.$
Here, the matrix $H(\theta^{k},\phi^{k},\lambda^{k})$ is defined as:
$H(\theta^{k},\phi^{k},\lambda^{k})=A^{-1}+\lambda^{k}A^{-1}BS^{-1}B^{T}A^{-1},$ (10)
where $S=C-\lambda^{k}B^{T}A^{-1}B$. The terms $A,B,$ and $C$ denote the derivatives of the Lagrangian $\mathcal{L}$:
$A$ $=\nabla^{2}_{\phi}\mathcal{L},\quad B=\nabla^{2}_{\phi\lambda}\mathcal{L},\quad C=\nabla_{\lambda}\mathcal{L},$ (11)
@@
$\mathcal{L}$ $=J(\theta,\phi)+\lambda\mathbb{E}_{\mathcal{D}_{\mu}}\left[\mathrm{KL}(P_{\bar{\phi}}\|P_{\phi})-\epsilon\right].$
@@
Eq. (9) updates the dual variable $\lambda$ for the constrained problem $\min_{\phi\in\Phi}J(\theta,\phi)$, solved alongside $\theta$ and $\phi$. Compared to Eq. (8), the policy updates integrate gradient from the best-response dual variable, which is a function of $\pi_{\theta}$. This allows the policy optimization to account for the implicit influence of constraint satisfaction, leading to a more informed learning process. The second line in Eq. (9) corresponds to a primal-dual method for solving $\min_{\phi\in\Phi}J(\theta,\phi)$, where $\cdot$$^{+}$ is a projection operation onto the non-negative real space. Widely used in constrained RL [^4] Viewing $P_{\phi}$ and $\pi_{\theta}$ as the ”policy” and ”world model” respectively, $\min_{\phi\in\Phi}J(\theta,\phi)$ is converted into a typical constrained RL problem., this method can achieve low suboptimality if the world model’s parameterization has sufficient representational capacity, as detailed in Paternain et al. (2023). Additionally, the first line in Eq. (9) represents a gradient ascent step corresponding to the objective function $\max_{\theta}\{J(\theta,\phi^{*}(\theta))\mid\phi^{*}(\theta)\in\arg\min_{\phi\in\Phi}J(\theta,\phi)\}$, of which the derivation is provided in Appendix F. According to the learning dynamics in Stackelberg games and constrained optimization (Fiez et al., 2020; Paternain et al., 2023), the learning rates in Eq. (9) should satisfy $\eta_{\phi}^{k}\gg\eta_{\lambda}^{k}\gg\eta_{\theta}^{k}$ for convergence.
@@ 5 Practical Algorithm: ROMBRL
For a discounted MDP with finite horizon $h$, $J(\theta,\phi)=\mathbb{E}_{\tau\sim P(\cdot;\theta,\phi)}\left[\sum_{j=0}^{h-1}\gamma^{j}r_{j}\right]$, $\tau=(s_{0},a_{0},r_{0},\cdots,s_{h})$ and $P(\tau;\theta,\phi)=d_{0}(s_{0})\prod_{j=0}^{h-1}\pi_{\theta}(a_{j}|s_{j})P_{\phi}(r_{j},s_{j+1}|s_{j},a_{j})$. To compute the first-order and second-order derivatives in Eqs. (7) - (9), we have the following theorem:
Theorem 5.1.
@@
$\nabla_{\theta}J(\theta,\phi)$ $=\mathbb{E}_{\tau\sim P(\cdot;\theta,\phi)}\left[\nabla_{\theta}\Psi(\tau,\theta)\right],$ (12)
$\nabla_{\phi}J(\theta,\phi)$ $=\mathbb{E}_{\tau\sim P(\cdot;\theta,\phi)}\left[\nabla_{\phi}\Psi(\tau,\phi)\right],$
@@
$\nabla_{\phi}^{2}J(\theta,\phi)$ $=\mathbb{E}_{\tau\sim P(\cdot;\theta,\phi)}\bigg[\nabla_{\phi}\Psi(\tau,\phi)\nabla_{\phi}\log P(\tau;\theta,\phi)^{T}$
$\quad+\nabla_{\phi}^{2}\Psi(\tau,\phi)\bigg].$
@@
In a practical algorithm, the expectation terms in Eqs. (12) and (56) are replaced with corresponding unbiased estimators (the sample means). Estimators of the first-order derivatives can be efficiently computed using automatic differentiation, whose space and time complexity scale linearly with the number of parameters in the models, namely $N_{\theta}$ and $N_{\phi}$. However, computing the second-order derivatives, i.e., $\nabla^{2}_{\phi}\Psi(\pi,\phi)$ in $\nabla^{2}_{\phi}J(\theta,\phi)$ and $\nabla^{2}_{\phi}\log P_{\phi}$ in $\nabla^{2}_{\phi}\mathcal{L}(\theta,\phi,\lambda)$, can be costly. Instead of substituting the second-order terms with $cI$ as in Wang et al. (2021), where $I$ is the identity matrix, we propose the following approximations:
$\mathbb{E}_{\tau\sim P(\cdot;\theta,\phi)}\left[\nabla_{\phi}^{2}\Psi(\tau,\phi)\right]$ (13)
$\quad\approx-\mathbb{E}_{\tau}\bigg[\sum_{i=0}^{h-1}\left(\sum_{j=i}^{h-1}\gamma^{j}r_{j}\right)\times F(s_{i},a_{i},r_{i},s_{i+1};\phi)\bigg],$
$\mathbb{E}_{(s,a,r,s^{\prime})\sim P_{\bar{\phi}}\circ\mathcal{D}_{\mu}(\cdot)}\left[\nabla^{2}_{\phi}\log P_{\phi}(r,s^{\prime}|s,a)\right]$
$\quad\approx-\mathbb{E}_{(s,a,r,s^{\prime})\sim P_{\bar{\phi}}\circ\mathcal{D}_{\mu}(\cdot)}\left[F(s,a,r,s^{\prime};\phi)\right],$
where $F(\cdot;\phi)=\nabla_{\phi}\log P_{\phi}\nabla_{\phi}\log P_{\phi}^{T}$ is the Fisher Information Matrix. Thus, we substitute $\nabla^{2}_{\phi}\log P_{\phi}$ with F-F, based on the fact (Amari, 2016) that $\mathbb{E}_{x\sim P_{\phi}(\cdot)}\left[\nabla^{2}_{\phi}\log P_{\phi}(x)+\nabla_{\phi}\log P_{\phi}(x)\nabla_{\phi}\log P_{\phi}(x)^{T}\right]=0$. In Appendix H, we discuss the approximation errors that arise when using Eq. (13).
The other computational bottleneck is computing the inverse matrix $A^{-1}$ in Eq. (9) [^5] Since $S$ in Eq. (9) is a scalar, $S^{-1}$ can be easily computed.. We choose not to use an identity matrix in place of $A^{-1}$ as done in Nikishin et al. (2022). As aforementioned, $A=\nabla^{2}_{\phi}\mathcal{L}(\theta,\phi,\lambda)$ is substituted with its sample-based estimator $\hat{A}$, which is defined as follows:
@@
where $U,V\in\mathbb{R}^{N_{\phi}\times m}$; $X,Y,Z\in\mathbb{R}^{N_{\phi}\times M}$; $m$ and $M$ represent the number of sampled trajectories and sampled transitions, respectively. Specifically, the $i$-th columns of $U$ and $V$ are given by $\nabla_{\phi}\Psi(\tau(i),\phi)/\sqrt{m}$ and $\nabla_{\phi}\log P(\tau(i);\theta,\phi)/\sqrt{m}$, respectively, where $\tau(i)$ denotes the $i$-th trajectory sampled from $P(\cdot;\theta,\phi)$. The $i$-th column of $Z$ is given by $\nabla_{\phi}\log P_{\phi}(r_{i},s^{\prime}_{i}|s_{i},a_{i})\cdot\sqrt{\lambda/M}$, based on a transition sampled from $P_{\bar{\phi}}\circ\mathcal{D}_{\mu}(\cdot)$. Additionally, by randomly sampling a trajectory $\tau$ and a time step $t\sim\text{Uniform}(0,h-1)$, we obtain the columns of $X$ and $Y$ as $\left(\sum_{j=t}^{h-1}\gamma^{j}r_{j}\right)\nabla_{\phi}\log P_{\phi}(s_{t+1},r_{t}|s_{t},a_{t})\cdot\sqrt{h/M}$ and $\nabla_{\phi}\log P_{\phi}(s_{t+1},r_{t}|s_{t},a_{t})\cdot\sqrt{h/M},$ respectively. Given that $m,M\ll N_{\theta},N_{\phi}$, each term in $\hat{A}$ is a low rank matrix and so we can apply Woodbury matrix identity (Fiez et al., 2020) to efficiently compute $\hat{A}^{-1}$.